OpenStudy (ttp):

Last week Brandon and Marcee paid a total of seventeen dollars to go to the movies. This week four of their friends are joining them. How much can they expect to pay altogether? thirty-four dollars fifty-one dollars sixty-eight dollars eighty-five dollars

OpenStudy (anonymous):

with interest.

OpenStudy (anonymous):

fifty one dollars

OpenStudy (anonymous):

wrong its 68 dollars

OpenStudy (anonymous):

how is it 68 ?

OpenStudy (anonymous):

Last week Brandon and Marcee paid a total of seventeen dollars to go to the movies. This week four of their friends are joining them. How much can they expect to pay altogether? 17 dollars for fisrt time then multiply that by the four friends cause it will stay the same price so 17x4=68

OpenStudy (anonymous):

first*

OpenStudy (anonymous):

Nope , your wrong . Were going to add . 17 for brandon and marcee then its going to be 34 for the 4 friends . add that up 17 + 34 = 51

OpenStudy (anonymous):

whatever you say. lol

OpenStudy (anonymous):

haha okay . @TTP its 51 :)

OpenStudy (anonymous):

@TTP it's 68

OpenStudy (mrnood):

@Nick123cross1 I think you have misread the question. The two guys paid a TOTAL of $17 - not $2 each. So 4 extra people would cost an extra 34 - total = $51

OpenStudy (anonymous):

exactly @Nick123cross1 its 51 :)

OpenStudy (anonymous):

ok

OpenStudy (texaschic101):

if it cost 2 people $17, then it is 8.50 per person, so 17 + 8.5(4) = 17 + 34 = 41

OpenStudy (texaschic101):

oops...51

OpenStudy (ttp):

Thank you @texaschic101 :)

OpenStudy (texaschic101):

no problem :)